A health food store sells seasoning blends in bulk. One blend contains 20% basil. Shelia wants to add pure basil to some 20% blend to make 16 ounces of her own 30% blend. Let b represent the amount of basil Shelia should add to the 20% blend.
Answer by checkley71(8403)

.20B+(16-B)=.3*16
.20B+16-B=4.8
-.8B=4.8-16
-.8B=11.2
B=-11.2/-.8
B=14 OZ. 20% BASIL.
THUS 16-14=2 OZ. OF PURE BASIL IS NEEDED.
PROOF
.2*14+2=.3*16
2.8+2=4.8
4.8=4.8